OpenStudy (anonymous):

How would you convert the repeating, non terminating decimal into a fraction? Explain the process as you solve the problem 0.1515

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

. First, count how many are not repeating: 0 Subtract the number from the whole decimal: 0.15-0=0.15 Now, put a 9 for every number repeating The answer is 15/99 Reduce: 5/33

OpenStudy (anonymous):

this good?

hero (hero):

Let x = 0.1515... 100x = 15.1515... 100x - x = 15 99x = 15 x = 15/99 x = 5/33

OpenStudy (anonymous):

ok so 100*0.1515

OpenStudy (anonymous):

yes

hero (hero):

But what were you going to say @amistre64 ?

OpenStudy (amistre64):

i was thinking of 15/100+15/10000+15/1000000+... and decided it was just not worth the effort to try to go on :)

hero (hero):

@amistre64, you never cease to humor me. I don't know whether you're just 'funnin' or not.

OpenStudy (amistre64):

a geometric sum was all i was considering ;)

hero (hero):

I just simply did was I was taught to do. I wasn't aware of any other methods to solve.

OpenStudy (amistre64):

there is a pattern:\[\frac{15}{100}+\frac{15}{100^2}+\frac{15}{100^3}+\frac{15}{100^4}+...\] \[15\left[(\frac{1}{100})+(\frac{1}{100})^2+(\frac{1}{100})^3+(\frac{1}{100})^4+...\right]\] \[\frac{15}{100}\left[1+(\frac{1}{100})+(\frac{1}{100})^2+(\frac{1}{100})^3+(\frac{1}{100})^4+...\right]\] this gives us a geometric series in 1/100 \[\frac{15}{100}\lim_{n\to \infty}\frac{1-.01^n}{1-.01}\] \[\frac{15}{100}\frac{1}{1-.01}\] \[\frac{15}{100}\frac{1}{.99}\] \[\frac{15}{99}...\]

OpenStudy (amistre64):

the formula for the sun of a geometric series is what you had already expressed is all

hero (hero):

Nice job presenting the alternative method @amistre64.